The sum of the interior angle measures of a convex polygon is 1620 degrees How many sides does it have?
Answer by stanbon(57307)

The sum of the interior angle measures of a convex polygon is 1620 degrees How many sides does it have?
----------------------------
Sum = (n-2)180
1620 = (n-2)180
n-2 = 9
n = 11 sides
=================
Cheers,
Stan H.
===============